OpenStudy (anonymous):

the demand for marazapan is 40 ounces per week when the price is $5 per ounce, but is only 35 ounces when the price rises to $7 per ounce. what is the demand as a function of price? (assuming a linear relationship between demand D and price per pound P)

OpenStudy (anonymous):

Let's say the demand goes on the y-axis, and the price goes on the x-axis. To get price per pound, multiply your prices per ounce by 16. To get demand in pounds, divide the demand in ounces by 16. Your points on the graph are: \begin{matrix}P_1 & D_1 & P_2 & D_2 \\ (80, & 2.5) & (112, & 2.1875)\end{matrix}You can use these to find the slope of the line that joins them with \[m=\frac{D_2-D_1}{P_2-P_1}\]And then you should use the equation of the line formula:\[D-D_1=m(P-P_1)\]